OpenStudy (anonymous):

indefinite integral of x sqrt of 2x+1 help!!

OpenStudy (anonymous):

chain rule

OpenStudy (anonymous):

derivative of x times sqrt of 2x+1 + derivative of sqrt of 2x+1 times x

OpenStudy (anonymous):

oh oops its integral LOL

OpenStudy (anonymous):

same thing

OpenStudy (anonymous):

yea lol @11

OpenStudy (anonymous):

you sure ?? @11calcBC

OpenStudy (anonymous):

no i mean the same rule apply to integration...lol

OpenStudy (anonymous):

im lost @11calcBC

OpenStudy (anonymous):

lol try substitution

OpenStudy (anonymous):

let u = 2x + 1

OpenStudy (anonymous):

ok I have all of that I just don't know wat goes after :o @11calcBC

OpenStudy (anonymous):

wuts ur resulting equation?

OpenStudy (anonymous):

with u instead of x

OpenStudy (anonymous):

after that it's just applying regular integration rules and ur done

OpenStudy (anonymous):

well I have u=2x+1 du=2dx and x=(u-1)/2 so I plug in what x= ??? @11calcBC

OpenStudy (anonymous):

well if u=2x+1, x would equal (u -1)/2 now wouldn't it?

OpenStudy (anonymous):

then multiple sqrt u with (u -1)/2, take out the 1/2 (its constant) and integrate

OpenStudy (anonymous):

oh! I think I get it! thanks! @11calcBC

OpenStudy (anonymous):

no problem mate good luck!

OpenStudy (anonymous):

thanks!! @11calcBC